# Scaling Behavior of Discrete Diffusion Language Models

### v1

@@ 4.2 Scaling Laws and Compute-Optimal Frontier
We find that both the optimal batch size and the optimal learning rate follow a very predictable trend (Fig. 3). The optimal batch size appears to depend primarily on the training horizon, with a remarkably strong, almost linear fit in the total number of training tokens. Similarly, the optimal learning rate follows a power law in the optimal batch size, where we assume that the batch size is chosen to be optimal for the given number of training tokens. Recent literature on scaling ALMs has reported similar predictable trends for both batch size and learning rate (bi2024deepseek; bergsma2025power). A more detailed analysis of the effect of model size and noise type on optimal hyperparametersis provided in Appendix A.4. Due to the predictability of the optimal learning rate, we sweep between only 2–3 different learning rates around the known optimal values for each batch size. Across all five noise types the resulting grid search spans 510 runs. See App. A.3 for details.
@@
hoffmann2022training 0.49 0.51
shuai2024scaling 0.464 0.536
bi2024deepseek 0.5243 0.4757
nie2025scaling 0.644 0.356 -0.0633
@@
The scaling laws describe the compute-optimal model size $M^{*}$ (in non-embedding FLOPs-per-token), training set size $D^{*}$ (in terms of tokens) and training loss $L^{*}$ (in terms of ELBO) as a function of training compute $C$ (in non-embedding FLOPs), following the methodology of bi2024deepseek. The scaling behavior reported by nie2025scaling and ni2025training differs considerably from ours, which could be due to differences in hyperparameters and/or training data. $2\sigma$-confidence intervals based on standard bootstrapping are given as subscripts.
To determine compute-optimal settings for each model size $M$ (in non-emb. FLOPs-per-token) and dataset size (in tokens), we select a set of target FLOPs and scan each observed loss curves for the loss value (in terms of the true ELBO, not the surrogate loss) achieved at the given target FLOPs. For smoothing, we apply a locally linear fit around the closest point to the target loss and determine the FLOP amount at which the target loss is crossed based on the linear fit. To fit the scaling laws, we adopt the approach based on iso-FLOP profiles from hoffmann2022training (Approach 2), as we find that the approach based on a parametric loss function (Approach 3) is unreliable and does not fit our data well. To approximate the number of FLOPs-per-token $M$ (as a measure of model expressivity), we follow bi2024deepseek and use $M=6P+12LDN$ (with non-emb. parameters $P$, layer count $L$, hidden size $D$, sequence length $N$) over to the more crude yet widely used $M=6P$ approximation.
Remarkably, we find a consistent trend in the scaling behavior of different noise types, with more uniform noise scaling more favorably with increased compute, requiring more parameters and less data to train compute-optimally. This is especially significant as the size of pre-training datasets is beginning to saturate while compute is continuing to become more abundant. Moreover, given that prior work has found comparable scaling behavior between autoregressive and masked diffusion models (nie2025scaling), this suggests that uniform diffusion models have the potential to outscale existing autoregressive training recipes. This finding is consistent with the notion that going from autoregressive modeling to masked diffusion to uniform diffusion imposes progressively less structure on the generation process and therefore less inductive bias, allowing it to scale more effortlessly with increased compute. Nevertheless, some limitations apply: Scaling coefficients can fluctuate across datasets depending on the data composition (bi2024deepseek), thus making our numbers not directly comparable with those of hoffmann2022training and shuai2024scaling.
@@ 4.4 Ablating the Effect of Learning Rate Annealing
For the sake of reducing the computational burden of scaling law estimation, we use a warmup-stable learning rate schedule without annealing, as outlined in Section 3. While we argue that this is a principled choice due to modern pre-training recipes spending most of the steps in the constant-LR regime and often treating the annealing as a separate phase (swissai2025apertus; team2025kimi; allal2025smollm2), we empirically investigate the the effect of learning rate annealing and find that it brings a constant improvement of 2.45% $\pm$ 0.138%88899% confidence interval, obtained through normal approximation. and does not affect optimal hyperparameters. Specifically, we investigate two settings: First, we fix the token budget while varying the batch size, sweeping the learning rate for each batch size both with and without annealing (Fig. 5(a)). This reveals that both learning rate and batch size have stable optima that are largely unaffected by annealing, which only shifts the final loss by a constant factor. Second, we investigate how the annealed performance evolves over the course of a single training run (Fig. 5(b)), again finding that the shape of the annealed loss closely matches the unannealed trajectory. Finally, we plot and extrapolate the correlation between the annealed and unannealed loss (Fig. 5(c)), revealing that the constant improvement continues to hold even for our scaled-up 3B and 10B parameter runs. [^9] The unannealed losses for the scaled-up runs are obtained through log-log extrapolation to the final 20% of steps based on the most recent steps before annealing begins. We therefore conclude that all scaling laws can safely be estimated without learning rate annealing, significantly reducing the number of ablation runs and compute requirements. The constant 2.45% factor can easily be incorporated at any point to account for the effect of adding annealing during the final training.
@@ 4.5 Relation between batch size and step count
At first sight, the mere existence of a positive $B_{\min}(L)$ and of a token-optimal batch size might appear to contradict classical results that advocate for small batches (keskar2016large; masters2018revisiting; smith2020generalization). However, those analyses typically assume a fixed dataset, arbitrarily many passes over the data, and focus on regularization effects of gradient noise which helps improve generalization. By contrast, our setting is internet-scale pre-training with an effective number of epochs below one and no observed overfitting. This difference in assumptions resolves the apparent tension with classical small-batch generalization results and is consistent with recent studies in ALM pre-training, which also report a token-dependent optimal batch size and diminishing returns beyond a critical batch size (hu2024minicpm; shuai2024scaling; bergsma2025power).
@@ A.1 Downstream Evaluations
We report the downstream performance according to a range of standard NLP benchmarks in Table 2. While the overall performance correlates with training ELBO, there appears to be a slight pattern of uniform diffusion performing comparatively better on reasoning-heavy tasks (ARC-E, ARC-C, GSM8k) and masked diffusion performing slightly better on knowledge-heavy tasks (PIQA, OBQA, BoolQ). The comparatively poor performance on GSM8k can be explained by the fact that our pre-training data is purely based on Nemotron-CC (su2024nemotron) and contains no dedicated math or coding data. This explanation is corroborated by the fact that all three checkpoints have a 0% pass rate on HumanEval (chen2021evaluating). The benchmarks were conducted using the lm-evaluation-harness [^10] https://github.com/EleutherAI/lm-evaluation-harness and the considered benchmarks are ARC-E, ARC-C (clark2018think), WinoGrande (sakaguchi2021winogrande), PIQA (bisk2020piqa), OpenBookQA (mihaylov2018can), BoolQ (clark2019boolq), and GSM8k (cobbe2021training). Except for GSM8k, these benchmarks are multiple-choice questions where the answer is selected based on likelihood. We report the best accuracy between 128/256 denoising steps and fill any unused context with tokens sampled from the prior distribution.
For GSM8k, we use a confidence-based sampling algorithm, similar to what is often done for masked diffusion (nie2025large; kim2025train). To this end, we extend confidence-based sampling to uniform diffusion and propose a generalized confidence heuristic that is applicable to both masked and uniform diffusion models. Specifically, in each step we fully denoise the top $k=1$ position that maximizes
@@ A.4 Optimal Hyperparameters
In Table 6 we report the scaling behavior of optimal hyperparameters (batch size and learning rate) of different noise types and model sizes individually in order to test whether optimal values depend on either of these quantities. We report the slope of a linear fit on a log-log scale along with its $R^{2}$ value as a measure of goodness of fit. While there are slight fluctuations in the scaling exponents grouped by model size, there is no consistent pattern for both optimal batch size and learning rate, with 99% confidence intervals often overlapping. More data would be needed to form a definitive answer, but based on the available observations we argue that the optimal batch size and learning rate are, more likely than not, unaffected by model size. Note that the independence of learning rate to model size is an expected consequence of using CompleteP (dey2025completeP). On the other hand, grouping by noise type reveals a consistent pattern of more uniform noise favoring larger batches. Nevertheless, the difference is rather small and unlikely to have practical implications as batch sizes are commonly rounded to a power of two for computational efficiency. The fits split by noise type are also visualized in Figure 7 (optimal batch size) and Figure 8 (optimal learning rate).
While the optimal batch size does not show any signs of saturation in our experiments, there likely exists a critical batch size above which increasing the size of the batch yields diminishing returns. The existence of such a batch size is well-established for autoregressive models (mccandlish2018empirical; shuai2024scaling) and is typically around $10^{6}$ tokens. Further, these findings may not generalize to multi-epoch training setups, as we operate in the sub-epoch training regime where overfitting is not a concern. Further research will be required to understand how optimal and critical batch sizes behave in multi-epoch training for both DLMs and ALMs.
@@
In Figure 9 we plot the optimal learning rates for each noise type, model size, batch size, and training horizon (in steps). For each batch size, there appears to be a power-law relations between optimal learning rate and training steps that is largely independent of model size, which is the expected consequence of using CompleteP (dey2025completeP). Furthermore, the optimal learning rate appears to increase with larger batches and decrease with larger training horizons, both of which are expected behaviors.
@@ Appendix B On the Difficulty of Uniform Diffusion
In Section 1, we argue that uniform diffusion is a strictly more difficult version of masked diffusion since assuming access to additional information about which tokens are noisy and noise-free reduces uniform diffusion to exactly masked diffusion. Therefore, removing access to this information can never make the task easier. A similar, more formal argument is raised by amin2025masking, saying that uniform diffusion has to learn the time of transition in addition to the transitions themselves, which is not the case for masked diffusion. The authors claim that this provides an inherent, theoretical advantage of masked diffusion over uniform diffusion, which at first glance may appear to contradict our results. However, we argue that both papers provide compatible findings and explanations: Since uniform diffusion is a strictly more difficult task, it requires more modeling capacity, resulting in a likelihood gap at small scales. Borrowing terminology from amin2025masking, small models struggle to learn both the transition schedule and the transitions themselves, but as we increase the size of the models, this becomes less of an issue since there is enough capacity to learn both. This can help further explain why uniform diffusion calls for scaling model size more quickly than data (compared to masked diffusion) and why the likelihood gap between masked and uniform diffusion shrinks with scale.
@@ Appendix D Scaling Coefficients
In this section we provide a complete report of our scaling law estimation methodology along with all fitted coefficients for the different variants. First, in addition to the more accurate FLOP/tok approximation proposed by bi2024deepseek, $M=6P+12LDN$, which we refer to as “Method 1”, we also report results for the simpler $M=6P$ approximation (kaplan2020scaling; hoffmann2022training), which is widely used in the literature (grattafiori2024llama; nie2025scaling; shuai2024scaling) and referred to as “Method 2”. We also report results both with and without smoothing applied to the iso-FLOP observations (referred to as “raw” and “sq. fit” respectively). It is standard to fit a parabola to observed iso-FLOP losses and taking the minimum thereof (grattafiori2024llama; bi2024deepseek) as a way of smoothing the observations, and while we find that both approaches qualitatively agree, the low resolution on the model sizes used results in a more brittle fit with broader confidence intervals for the “raw” data. Further, we also ablate whether or not to include an irreducible term in the power law, i.e. whether to use $f(C)=AC^{\alpha}$ or $f(C)=AC^{\alpha}+E$ as a hypothesis, and find that, especially for the smoothed data, the irreducible term is typically very small or zero. We therefore conclude that the irreducible term is too small to accurately model, and that omitting it from the hypothesis results in a more robust and comparable fit.
@@
Figure 10: The fitted scaling coefficients differ systematically between FLOP estimation techniques: Method 1 uses the FLOP estimation technique proposed by bi2024deepseek whereas method 2 uses the classic approach by hoffmann2022training. Furthermore, fitting on interpolated data (squared fit) produces tighter confidence bounds and better scaling exponents. Shaded regions denote 95% confidence intervals obtained via standard bootstrapping on the aggregated data points.
Table 7: Compute-constrained scaling coefficients for all noise types, metrics, and methodologies, obtained by fitting the power law $AC^{\alpha}$ (where $C=MD$) to the observed data. Method 1 uses the FLOP/tok estimation from bi2024deepseek while method 2 uses the classic $M=6P$ approximation (hoffmann2022training). ‘raw’ interpolation refers to taking the optimal observed value for a given iso-FLOP target (i.e. no interpolation), whereas the ‘sq. fit’ data is obtained by fitting a parabola to the observed values and taking the optimum thereof. Smallest scaling coefficients are bolded.
@@
Figure 11: Compute-optimal scaling laws fitted to interpolated values (squared fit). The FLOP estimation methodologies by bi2024deepseek (Method 1) and hoffmann2022training (Method 2) differ significantly since the FLOP-approximation used by hoffmann2022training ($M=6P$) systematically underestimates the total number of FLOPs executed during training.
Figure 12: Compute-optimal scaling laws fitted on non-interpolated observations. Compared to interpolated values, the optimal observed values can deviate significantly from the compute-optimal trend due to the small number of unique model sizes used in the sweep. Method 1 uses the FLOP estimation technique from bi2024deepseek, whereas Method 2 follows hoffmann2022training.

### v2

@@ 4.2 Scaling Laws and Compute-Optimal Frontier
We find that both the optimal batch size and the optimal learning rate follow a very predictable trend (Fig. 3). The optimal batch size appears to depend primarily on the training horizon, with a remarkably strong, almost linear fit in the total number of training tokens. Similarly, the optimal learning rate follows a power law in the optimal batch size, where we assume that the batch size is chosen to be optimal for the given number of training tokens. Recent literature on scaling ALMs has reported similar predictable trends for both batch size and learning rate (Bi et al., 2024; Bergsma et al., 2025). A more detailed analysis of the effect of model size and noise type on optimal hyperparametersis provided in Appendix A.4. Due to the predictability of the optimal learning rate, we sweep between only 2–3 different learning rates around the known optimal values for each batch size. Across all five noise types the resulting grid search spans 510 runs. See App. A.3 for details.
@@
Hoffmann et al. (2022) 0.49 0.51
Shuai et al. (2024) 0.464 0.536
Bi et al. (2024) 0.5243 0.4757
Nie et al. (2025a) 0.644 0.356 -0.0633
@@
The scaling laws describe the compute-optimal model size $M^{*}$ (in non-embedding FLOPs-per-token), training set size $D^{*}$ (in terms of tokens) and training loss $L^{*}$ (in terms of ELBO) as a function of training compute $C$ (in non-embedding FLOPs), following the methodology of Bi et al. (2024). The scaling behavior reported by Nie et al. (2025a) and Ni et al. (2025) differs considerably from ours, which could be due to differences in hyperparameters and/or training data. $2\sigma$-confidence intervals based on standard bootstrapping are given as subscripts.
To determine compute-optimal settings for each model size $M$ (in non-emb. FLOPs-per-token) and dataset size (in tokens), we select a set of target FLOPs and scan each observed loss curves for the loss value (in terms of the true ELBO, not the surrogate loss) achieved at the given target FLOPs. For smoothing, we apply a locally linear fit around the closest point to the target loss and determine the FLOP amount at which the target loss is crossed based on the linear fit. To fit the scaling laws, we adopt the approach based on iso-FLOP profiles from Hoffmann et al. (2022) (Approach 2), as we find that the approach based on a parametric loss function (Approach 3) is unreliable and does not fit our data well. To approximate the number of FLOPs-per-token $M$ (as a measure of model expressivity), we follow Bi et al. (2024) and use $M=6P+12LDN$ (with non-emb. parameters $P$, layer count $L$, hidden size $D$, sequence length $N$) over to the more crude yet widely used $M=6P$ approximation.
Remarkably, we find a consistent trend in the scaling behavior of different noise types, with more uniform noise scaling more favorably with increased compute, requiring more parameters and less data to train compute-optimally. This is especially significant as the size of pre-training datasets is beginning to saturate while compute is continuing to become more abundant. Moreover, given that prior work has found comparable scaling behavior between autoregressive and masked diffusion models (Nie et al., 2025a), this suggests that uniform diffusion models have the potential to outscale existing autoregressive training recipes. This finding is consistent with the notion that going from autoregressive modeling to masked diffusion to uniform diffusion imposes progressively less structure on the generation process and therefore less inductive bias, allowing it to scale more effortlessly with increased compute. Nevertheless, some limitations apply: Scaling coefficients can fluctuate across datasets depending on the data composition (Bi et al., 2024), thus making our numbers not directly comparable with those of Hoffmann et al. (2022) and Shuai et al. (2024).
@@ 4.4 Ablating the Effect of Learning Rate Annealing
For the sake of reducing the computational burden of scaling law estimation, we use a warmup-stable learning rate schedule without annealing, as outlined in Section 3. While we argue that this is a principled choice due to modern pre-training recipes spending most of the steps in the constant-LR regime and often treating the annealing as a separate phase (Project Apertus, 2025; Kimi Team, 2025; Allal et al., 2025), we empirically investigate the the effect of learning rate annealing and find that it brings a constant improvement of 2.45% $\pm$ 0.138%88899% confidence interval, obtained through normal approximation. and does not affect optimal hyperparameters. Specifically, we investigate two settings: First, we fix the token budget while varying the batch size, sweeping the learning rate for each batch size both with and without annealing (Fig. 5(a)). This reveals that both learning rate and batch size have stable optima that are largely unaffected by annealing, which only shifts the final loss by a constant factor. Second, we investigate how the annealed performance evolves over the course of a single training run (Fig. 5(b)), again finding that the shape of the annealed loss closely matches the unannealed trajectory. Finally, we plot and extrapolate the correlation between the annealed and unannealed loss (Fig. 5(c)), revealing that the constant improvement continues to hold even for our scaled-up 3B and 10B parameter runs. [^9] The unannealed losses for the scaled-up runs are obtained through log-log extrapolation to the final 20% of steps based on the most recent steps before annealing begins. We therefore conclude that all scaling laws can safely be estimated without learning rate annealing, significantly reducing the number of ablation runs and compute requirements. The constant 2.45% factor can easily be incorporated at any point to account for the effect of adding annealing during the final training.
@@ 4.5 Relation between batch size and step count
At first sight, the mere existence of a positive $B_{\min}(L)$ and of a token-optimal batch size might appear to contradict classical results that advocate for small batches (Keskar et al., 2016; Masters & Luschi, 2018; Smith et al., 2020). However, those analyses typically assume a fixed dataset, arbitrarily many passes over the data, and focus on regularization effects of gradient noise which helps improve generalization. By contrast, our setting is internet-scale pre-training with an effective number of epochs below one and no observed overfitting. This difference in assumptions resolves the apparent tension with classical small-batch generalization results and is consistent with recent studies in ALM pre-training, which also report a token-dependent optimal batch size and diminishing returns beyond a critical batch size (Hu et al., 2024; Shuai et al., 2024; Bergsma et al., 2025).
@@ 5 Conclusion
Acknowledgments
This research was supported with Cloud TPUs from Google’s TPU Research Cloud (TRC). We thank Erfan Zare Chavoshi for the development of and technical support related to the EasyDeL framework (Zare Chavoshi, 2023), which was used for training our models. Antonio Orvieto and Bernhard Schölkopf acknowledge the financial support of the Hector Foundation.
@@ A.1 Downstream Evaluations
We report the downstream performance according to a range of standard NLP benchmarks in Table 2. While the overall performance correlates with training ELBO, there appears to be a slight pattern of uniform diffusion performing comparatively better on reasoning-heavy tasks (ARC-E, ARC-C, GSM8k) and masked diffusion performing slightly better on knowledge-heavy tasks (PIQA, OBQA, BoolQ). The comparatively poor performance on GSM8k can be explained by the fact that our pre-training data is purely based on Nemotron-CC (Su et al., 2024) and contains no dedicated math or coding data. This explanation is corroborated by the fact that all three checkpoints have a 0% pass rate on HumanEval (Chen, 2021). The benchmarks were conducted using the lm-evaluation-harness [^10] https://github.com/EleutherAI/lm-evaluation-harness and the considered benchmarks are ARC-E, ARC-C (Clark et al., 2018), WinoGrande (Sakaguchi et al., 2021), PIQA (Bisk et al., 2020), OpenBookQA (Mihaylov et al., 2018), BoolQ (Clark et al., 2019), and GSM8k (Cobbe et al., 2021). Except for GSM8k, these benchmarks are multiple-choice questions where the answer is selected based on likelihood. We report the best accuracy between 128/256 denoising steps and fill any unused context with tokens sampled from the prior distribution.
For GSM8k, we use a confidence-based sampling algorithm, similar to what is often done for masked diffusion (Nie et al., 2025b; Kim et al., 2025). To this end, we extend confidence-based sampling to uniform diffusion and propose a generalized confidence heuristic that is applicable to both masked and uniform diffusion models. Specifically, in each step we fully denoise the top $k=1$ position that maximizes
@@ A.4 Optimal Hyperparameters
In Table 6 we report the scaling behavior of optimal hyperparameters (batch size and learning rate) of different noise types and model sizes individually in order to test whether optimal values depend on either of these quantities. We report the slope of a linear fit on a log-log scale along with its $R^{2}$ value as a measure of goodness of fit. While there are slight fluctuations in the scaling exponents grouped by model size, there is no consistent pattern for both optimal batch size and learning rate, with 99% confidence intervals often overlapping. More data would be needed to form a definitive answer, but based on the available observations we argue that the optimal batch size and learning rate are, more likely than not, unaffected by model size. Note that the independence of learning rate to model size is an expected consequence of using CompleteP (Dey et al., 2025). On the other hand, grouping by noise type reveals a consistent pattern of more uniform noise favoring larger batches. Nevertheless, the difference is rather small and unlikely to have practical implications as batch sizes are commonly rounded to a power of two for computational efficiency. The fits split by noise type are also visualized in Figure 7 (optimal batch size) and Figure 8 (optimal learning rate).
While the optimal batch size does not show any signs of saturation in our experiments, there likely exists a critical batch size above which increasing the size of the batch yields diminishing returns. The existence of such a batch size is well-established for autoregressive models (McCandlish et al., 2018; Shuai et al., 2024) and is typically around $10^{6}$ tokens. Further, these findings may not generalize to multi-epoch training setups, as we operate in the sub-epoch training regime where overfitting is not a concern. Further research will be required to understand how optimal and critical batch sizes behave in multi-epoch training for both DLMs and ALMs.
@@
In Figure 9 we plot the optimal learning rates for each noise type, model size, batch size, and training horizon (in steps). For each batch size, there appears to be a power-law relations between optimal learning rate and training steps that is largely independent of model size, which is the expected consequence of using CompleteP (Dey et al., 2025). Furthermore, the optimal learning rate appears to increase with larger batches and decrease with larger training horizons, both of which are expected behaviors.
@@ Appendix B On the Difficulty of Uniform Diffusion
In Section 1, we argue that uniform diffusion is a strictly more difficult version of masked diffusion since assuming access to additional information about which tokens are noisy and noise-free reduces uniform diffusion to exactly masked diffusion. Therefore, removing access to this information can never make the task easier. A similar, more formal argument is raised by Amin et al. (2025), saying that uniform diffusion has to learn the time of transition in addition to the transitions themselves, which is not the case for masked diffusion. The authors claim that this provides an inherent, theoretical advantage of masked diffusion over uniform diffusion, which at first glance may appear to contradict our results. However, we argue that both papers provide compatible findings and explanations: Since uniform diffusion is a strictly more difficult task, it requires more modeling capacity, resulting in a likelihood gap at small scales. Borrowing terminology from Amin et al. (2025), small models struggle to learn both the transition schedule and the transitions themselves, but as we increase the size of the models, this becomes less of an issue since there is enough capacity to learn both. This can help further explain why uniform diffusion calls for scaling model size more quickly than data (compared to masked diffusion) and why the likelihood gap between masked and uniform diffusion shrinks with scale.
@@ Appendix D Scaling Coefficients
In this section we provide a complete report of our scaling law estimation methodology along with all fitted coefficients for the different variants. First, in addition to the more accurate FLOP/tok approximation proposed by Bi et al. (2024), $M=6P+12LDN$, which we refer to as “Method 1”, we also report results for the simpler $M=6P$ approximation (Kaplan et al., 2020; Hoffmann et al., 2022), which is widely used in the literature (Grattafiori et al., 2024; Nie et al., 2025a; Shuai et al., 2024) and referred to as “Method 2”. We also report results both with and without smoothing applied to the iso-FLOP observations (referred to as “raw” and “sq. fit” respectively). It is standard to fit a parabola to observed iso-FLOP losses and taking the minimum thereof (Grattafiori et al., 2024; Bi et al., 2024) as a way of smoothing the observations, and while we find that both approaches qualitatively agree, the low resolution on the model sizes used results in a more brittle fit with broader confidence intervals for the “raw” data. Further, we also ablate whether or not to include an irreducible term in the power law, i.e. whether to use $f(C)=AC^{\alpha}$ or $f(C)=AC^{\alpha}+E$ as a hypothesis, and find that, especially for the smoothed data, the irreducible term is typically very small or zero. We therefore conclude that the irreducible term is too small to accurately model, and that omitting it from the hypothesis results in a more robust and comparable fit.
@@
Figure 10: The fitted scaling coefficients differ systematically between FLOP estimation techniques: Method 1 uses the FLOP estimation technique proposed by Bi et al. (2024) whereas method 2 uses the classic approach by Hoffmann et al. (2022). Furthermore, fitting on interpolated data (squared fit) produces tighter confidence bounds and better scaling exponents. Shaded regions denote 95% confidence intervals obtained via standard bootstrapping on the aggregated data points.
Table 7: Compute-constrained scaling coefficients for all noise types, metrics, and methodologies, obtained by fitting the power law $AC^{\alpha}$ (where $C=MD$) to the observed data. Method 1 uses the FLOP/tok estimation from Bi et al. (2024) while method 2 uses the classic $M=6P$ approximation (Hoffmann et al., 2022). ‘raw’ interpolation refers to taking the optimal observed value for a given iso-FLOP target (i.e. no interpolation), whereas the ‘sq. fit’ data is obtained by fitting a parabola to the observed values and taking the optimum thereof. Smallest scaling coefficients are bolded.
@@
Figure 11: Compute-optimal scaling laws fitted to interpolated values (squared fit). The FLOP estimation methodologies by Bi et al. (2024) (Method 1) and Hoffmann et al. (2022) (Method 2) differ significantly since the FLOP-approximation used by Hoffmann et al. (2022) ($M=6P$) systematically underestimates the total number of FLOPs executed during training.
Figure 12: Compute-optimal scaling laws fitted on non-interpolated observations. Compared to interpolated values, the optimal observed values can deviate significantly from the compute-optimal trend due to the small number of unique model sizes used in the sweep. Method 1 uses the FLOP estimation technique from Bi et al. (2024), whereas Method 2 follows Hoffmann et al. (2022).